OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 4. Find x when y = 8. Show all work.

OpenStudy (johnweldon1993):

y=k/3x when x = 6, y = 4 4=k/3*6 solve for k k=4*3*6=72 now when y = 8 Then 8=72/3x divide by 3 8=24/x x=24/8 x = 3